OpenStudy (anonymous):

what are the solutions for the quadratic equation f(x)=15x^2+x-2

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

\[-1+ or - \sqrt{41}/10\]

OpenStudy (anonymous):

How did you get that

OpenStudy (anonymous):

The Discriminant y=5x2+1x−2 discriminant =b2−4(a)(c)=12−4(5)(−2)=41 ( 2 real solutions ) The Work −b±b2−4(a)(c)‾‾‾‾‾‾‾‾‾‾‾√2(a)−1±12−4(5)(−2)‾‾‾‾‾‾‾‾‾‾‾‾‾√2(5)−1±41‾‾‾√10 The Actual Solutions x=0.5403124237432848 x=−0.7403124237432849

OpenStudy (anonymous):

it wont let me make the right signs hold on.

OpenStudy (anonymous):

okay